OpenStudy (anonymous):

If F(x) = 2x and G(x) = log x, find F [G(x)]

OpenStudy (anonymous):

plug g(x) instead of x in F(x) F(G(X))=2(logx)

OpenStudy (anonymous):

F(x) = 2(G(x)) = 2(log x)

OpenStudy (unklerhaukus):

\[=\log x^2\]